# Multi-modal Auto-regressive Modeling via Visual Words

### v1

@@ Abstract
Large Language Models (LLMs), benefiting from the auto-regressive modelling approach performed on massive unannotated texts corpora, demonstrates powerful perceptual and reasoning capabilities. However, as for extending auto-regressive modelling to multi-modal scenarios to build Large Multi-modal Models (LMMs), there lies a great difficulty that the image information is processed in the LMM as continuous visual embeddings, which cannot obtain discrete supervised labels for classification. In this paper, we successfully perform multi-modal auto-regressive modeling with a unified objective for the first time. Specifically, we propose the concept of visual words, which maps the visual features to probability distributions over LLM’s vocabulary, providing supervision information for visual modelling. We further explore the distribution of visual features in the semantic space within LMM and the possibility of using text embeddings to represent visual information. Experimental results and ablation studies on 5 VQA tasks and 4 benchmark toolkits validate the powerful performance of our proposed approach. Code are available at https://github.com/pengts/VW-LMM
Multi-modal Auto-regressive Modeling via Visual Words
[…]
1 Introduction
Figure 1: Performance of our proposed VW-LMM.
@@ 1 Introduction
Figure 2: The comparisons between different LMMs. (a) Mainstream methods focuses on predicting language responses in multi-modal contextual sequences that depend on visual information, where the visual information merely acts as contextual cues and does not serve as supervision. (b) Visual regression method takes regression task to predict the value of next visual feature and performs joint training with text. (c) Our multi-modal auto-regressive method use visual words to construct visual supervision labels and enabling multi-modal auto-regressive modelling with unified classification objective.
@@
Over the past year, Large Language Models (LLMs) have made impressive breakthroughs and successfully use language as a common interface for a wide variety of real-world tasks. Benefiting from the auto-regressive modelling approach performed on massive unannotated texts, LLM is able to learn general-purpose semantic information and powerful reasoning capabilities from natural language corpora. The success of LLMs attracts researchers to explore Large Multi-modal Models (LMMs), which aim to extend the powerful text-only perceptual and reasoning capabilities of LLMs to scenarios dealing with multi-modal inputs. However, as for extending auto-regressive modelling to multi-modal scenarios, there lies a great difficulty that the image information is processed in the LMM as continuous visual embeddings, which cannot obtain discrete supervised labels for classification. As a compromise solution, mainstream LMMs choose to compute losses only for the language portion of multi-modal interleaved sequence. As illustrated in figure  2(a), the training objective of mainstream methods Dai et al. (2023); Li et al. (2023a); Liu et al. (2023a); Lu et al. (2023); Bai et al. (2023a) focuses on predicting language responses in multi-modal contextual sequences that depend on visual information, where the visual information merely acts as contextual cues and does not serve as supervision. This unfair treatment of different modal information in multi-modal sequences lacks the process of learning different modal information utilising the inference capabilities of the LLM, severely limiting the potential of LMM and resulting in under-utilisation of the training data. Although recent works  Sun et al. (2023a, b) propose to unlock the LLM for text pre-training by using a regression task to predict the value of next visual feature in the pre-training phase (Fig.  2(b)), the inconsistent optimisation goals of its visual and linguistic components are not conducive to unified multi-modal auto-regressive modelling.  Bai et al. (2023b) has also proposed learning visual features using an auto-regressive classification tasks, but it uses a pre-trained image tokenizer, such as VQVAE or VQGAN, to cluster image features into a grid of discrete tokens which does not combine discrete visually supervised information with LLM.
@@
In this paper, we successfully perform multi-modal auto-regressive modeling with a unified objective for the first time. Specifically, we introduce the concept of visual words to construct representations of visual features in the language semantic space inside the LMM, thus implementing visual modelling of LMM in the form of classification rather than regression. We also propose the Visual Word guided Large Multi-modal Model (VW-LMM), a novel multi-modal model that inherits the successful learning paradigm of LLM in the pre-training task, i.e., predicting the next image/text token in an auto-regressive manner. With the help of visual words, it’s possible to train LMM with auto-regressive modeling without any specific architectural modifications, as shown in figure  2(c). Just as the LM head component is used in LLM to accomplish the mapping of text features to interpretable semantics, we correspondingly construct the VM head component, using the same structure, intent on obtaining the representation of each visual embedding in the language semantic space of LLM. For each visual embedding, we use the VM head to map it into probability distributions over the pre-trained vocabulary, which we call visual words. After that, the visual words corresponding to the image modal and the one hot label corresponding to the text modal are intertwined to form the supervised information for multi-modal auto-regressive modelling. Further, the pre-trained embeddings weights of LLM can be approximated as a set of complete bases of the language semantic space, i.e., the semantic information covered by the embeddings can basically cover the whole semantic space. Therefore, we further explored whether pseudo image features constructed with visual words and LLM’s pre-trained embeddings can convey visual information to the model. We conducted extensive experiments on five commonly used visual question answering benchmarks and four LMM-evaluating benchmark toolkits, and the experimental results demonstrate that our VW-LMM, by constructing visual words to introduce visual supervisory information, achieves the best performance among models of the same scale, and obtains vision-language understanding capability competitive to or even surpassing that of 13B or even larger scale models with a scale of 7B. Our main contributions are as follows:
@@
  • We propose the concept of visual words, which maps visual features into language semantic space, enabling LMM to perform auto-regressive modelling over multi-modal sequence.
@@
2 Method
Figure 3: The overview of our method. (a) The overall framework of the model. VW-LMM uses the VM head to transform visual features in multi-modal input sequences to probability distributions over LLM’s vocabulary (so-called visual words) to participate in visual modelling (b) Constructing pseudo image features with pre-trained embedding of LLM and visual words. (c) Demonstration of semantically closest tokens of each image patch in LMM.
2.1 Multi-modal Learning
To extend the powerful text-only perceptual and reasoning capabilities of LLM to scenarios dealing with multi-modal inputs, existing multi-modal learning method of LMM typically use the adapter structure to transform the visual features encoded by pre-trained visual backbone into the semantic space of LLM and construct multi-modal input sequences together with text embeddings. The adapter could be well-designed visual resampler Bai et al. (2023a); Dai et al. (2023), Multi-layer Perceptron (MLP) Liu et al. (2023a) or even simple Linear layer Liu et al. (2023b).
@@ 2.1 Multi-modal Learning
$X_{\mathrm{image}}=\mathrm{AD}(\mathrm{VE}(\mathrm{image})),$ (1)
@@
Subsequently, $X_{\mathrm{input}}$ is fed into the LLM for auto-regressive decoding. Assuming that the LLM mainly contains two important components: a language decoder $\Gamma$ and an LM head, the loss of classical multi-modal learning can be expressed as
$Loss_{\mathrm{LM}}$ $=\frac{1}{|S_{\mathrm{{LM}}}|}\sum_{n\in S_{\mathrm{{LM}}}}\left(-\sum_{i=0}^{C-1}P_{n}(i)\log Q_{n}(i)\right),$ (2)
@@
where $W_{\mathrm{LM}}$ is the optimizable parameters of $\mathrm{LMhead}$, set $S_{\mathrm{{LM}}}$ is the set of expected text output index of model and $C$ is the vocabulary size of model.
However, classical multi-modal learning neglects the supervision function of visual information on multi-modal auto-regressive modelling, which is still multi-modal based language modelling in essence. Our work proposes to leverage the great reasoning potential of LLM to facilitate both language modelling and visual modelling of LMM. To achieve this goal, the key component is to construct formally unified multi-modal generative objectives so that the model can be trained with auto-regressive manner. As shown in Figure 3(a), the architecture of VW-LMM consists of five components: a Visual encoder, a multi-modal decoder, an adapter for projection between visual modal and language modal, a Multi-modal Modeling head (MM head) for multi-modal modelling and a corresponding VM head for visual modelling. The embedding module can be regarded as part of multi-modal decoder. We initialise the multi-modal decoder and MM head in VW-LMM using the pre-trained LLM and its LM head. In contrast to classical multi-modal learning, VW-LMM uses VMhead to construct visual words for visual features as supervisory information, thus enabling the model to perform multi-modal auto-regressive modeling over the entire sequence.
2.2 Visual Words
@@
$v_{i}=\arg\min_{j}\operatorname{Cosine}(x_{i},e_{j}),\quad v_{i}\in[0,C-1],$ (3)
@@
We show the semantically closest tokens of image regions in figure 3(c), it can be observed that for the LMM with multi-modal alignment training, there is a certain correlation between the information contained in each vector of the image features and the actual content of the image. Moreover, this correlation can be explicitly expressed by the vocabulary of the LMM. This finding, on the one hand, proves that the visual features within the LMM are in a similar semantic space as the text embedding, and on the other hand, it provides a new perspective for the interpretability analysis of the LMM. Furthermore, existing LLM methods have shown that the hidden states of text features can be mapped to the model’s vocabulary through linear projection, enabling the extraction of interpretable semantics. Therefore, we propose using the linear projection to also map the visual features to the probability distribution over the model’s vocabulary, which we called visual words, to further strengthen the correlation between visual features and text embeddings.
@@
$Loss_{\mathrm{LM}}$ $=\frac{1}{|S_{\mathrm{{LM}}}|}\sum_{n\in S_{\mathrm{{LM}}}}\left(-\sum_{i=0}^{C-1}P_{n}(i)\log Q_{n}(i)\right),$ (4)
@@
The visual words of visual features can be represented as
$P^{\prime}$ $=\operatorname{Softmax}(W_{\mathrm{VM}}X_{\mathrm{image}}),$ (5)
@@
$Loss_{\mathrm{VM}}$ $=\frac{1}{|S_{\mathrm{{VM}}}|}\sum_{n\in S_{\mathrm{{VM}}}}\left(\sum_{i=1}^{C}P^{\prime}_{n}(i)\frac{\log P^{\prime}_{n}(i)}{\log Q_{n}(i)}\right).$ (6)
where set $S_{\mathrm{{VM}}}$ is the set of visual output index of model. It should be noted that in the training process of VW-LMM, the optimisation phase of VM head and multi-modal decoder are separated, and thus do not cause instability during training.
@@
$Loss_{\mathrm{MM}}=Loss_{\mathrm{LM}}+Loss_{\mathrm{VM}}.$ (7)
@@
2.3 Fused Text Embeddings as Pseudo Image Features
The text embeddings in the multi-modal decoder can be regarded as a set of base vectors within its semantic space. Therefore, we propose utilizing the visual words and text embeddings to construct pseudo image features, aiming to further explore the manifestation of visual features in the semantic space of the LMM.
@@ 2.3 Fused Text Embeddings as Pseudo Image Features
$X_{\mathrm{p}}=W_{\mathrm{embeddings}}\odot\operatorname{Softmax}\left((W_{\mathrm{VM}}X_{\mathrm{image}})\right)$ (8)
@@ 3 Experiments
IDEFICS-80B IDEFICS (2023) LLaMA-65B 224 60.0 45.2 36.0 30.9 54.5 38.1
InstructBLIP Dai et al. (2023) Vicuna-13B 224 49.5 33.4 63.1 50.7 78.9 25.6
BLIP-2 Li et al. (2023a) Vicuna-13B 224 41.0 41.0 19.6 61.0 42.5 85.3 22.4
LLaVA-v1.5 Liu et al. (2023a) Vicuna-13B 336 $*{}^{*}$ $*{}^{*}$ 53.6 71.6 61.3 85.9 67.7 63.6 35.4
InstructBLIP Dai et al. (2023) Vicuna-7B 224 49.2 34.5 60.5 50.1 36 23.7 26.2
IDEFICS-9B IDEFICS (2023) LLaMA-7B 224 50.9 38.4 35.5 25.9 48.2 25.2
Qwen-VL Bai et al. (2023a) Qwen-7B 448 $*{}^{*}$ $*{}^{*}$ 35.2 67.1 63.8 38.2 7.4
Qwen-VL-Chat Bai et al. (2023a) Qwen-7B 448 $*{}^{*}$ $*{}^{*}$ 38.9 68.2 61.5 60.6 56.7
LLaVA-v1.5 Liu et al. (2023a) Vicuna-7B 336 $*{}^{*}$ $*{}^{*}$ 50.0 66.8 58.2 85.9 64.3 58.3 30.5
MoE-LLaVA-2.7B×4-Top2 Lin et al. (2024) Phi-2-2.7B 336 $*{}^{*}$ $*{}^{*}$ 43.9 68.5 51.4 86.3 65.2 34.3
@@
Emu2-Chat Sun et al. (2023a) LLaMA-33B 448 $*{}^{*}$ $*{}^{*}$ 54.9 $*{}^{*}$ $*{}^{*}$ 48.5
Emu-I Sun et al. (2023b) LLaMA-13B 224 62.0 46.0 38.3 36.3
MM-Interleaved-SFT Tian et al. (2024) Vicuna-13B 224 $*{}^{*}$ $*{}^{*}$ 54.9 61.0
Unified-IO 2 Lu et al. (2023) UIO-2-6.8B 384 $*{}^{*}$ $*{}^{*}$ 87.7 71.5
DreamLLM Dong et al. (2023) Vicuna-7B 224 56.6 38.1 34.9
VL-GPT-I Zhu et al. (2023) LLaMA-7B 224 67.2 51.5 38.9
LaVIT-v2 Jin et al. (2023) LLaMA2-7B 224 68.3 47.9 41.0
VW-LMM Vicuna-7B 336 $*{}^{*}$ $*{}^{*}$ 48.3 68.1 57.6 85.9 65.9 59.8 31.3
VW-LMM Mistral-7B 336 $*{}^{*}$ $*{}^{*}$ 58.5 75.9 63.1 87.0 80.6 79.0 44.0

### v2

# Multi-modal Auto-regressive Modeling via Visual Tokens

@@ Abstract.
Large Language Models (LLMs), benefiting from the auto-regressive modelling approach performed on massive unannotated texts corpora, demonstrates powerful perceptual and reasoning capabilities. However, as for extending auto-regressive modelling to multi-modal scenarios to build Large Multi-modal Models (LMMs), there lies a great difficulty that the image information is processed in the LMM as continuous visual embeddings, which cannot obtain discrete supervised labels for classification. In this paper, we successfully perform multi-modal auto-regressive modeling with a unified objective for the first time. Specifically, we propose the concept of visual tokens, which maps the visual features to probability distributions over LLM’s vocabulary, providing supervision information for visual modelling. We further explore the distribution of visual features in the semantic space within LMM and the possibility of using text embeddings to represent visual information. Experimental results and ablation studies on 5 VQA tasks and 4 benchmark toolkits validate the powerful performance of our proposed approach. The code is released at https://github.com/pengts/VW-LMM.
Multimedia Foundation Models, Multi-modal Auto-regressive Modeling, Vision and Language, Modal Interpretation
copyright: acmlicensedjournalyear: 2024conference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10doi: 10.1145/3664647.3681685ccs: Computing methodologies Natural language generationccs: Computing methodologies Image representations
1. Introduction
Figure 1. Performance of our proposed VT-LMM. We set the unreported results in the original paper to 0 to avoid confusion.
@@ 1. Introduction
Figure 2. The comparisons between different LMMs. (a) Visual Context Method focuses on predicting language responses in multi-modal contextual sequences that depend on visual information, where the visual information merely acts as contextual cues and does not serve as supervision. (b) Visual regression method takes regression task to predict the value of next visual feature and performs joint training with text. (c) Our multi-modal auto-regressive method use visual tokens to construct visual supervision labels and enabling multi-modal auto-regressive modelling with unified classification objective.
@@
Over the past year, Large Language Models (LLMs) have made impressive breakthroughs and successfully use language as a common interface for a wide variety of real-world tasks. Benefiting from the auto-regressive modelling approach performed on massive unannotated texts, LLM is able to learn general-purpose semantic information and powerful reasoning capabilities from natural language corpora. The success of LLMs attracts researchers to explore Large Multi-modal Models (LMMs), which aim to extend the powerful text-only perceptual and reasoning capabilities of LLMs to scenarios dealing with multi-modal inputs. However, as for extending auto-regressive modelling to multi-modal scenarios, there lies a great difficulty that the image information is processed in the LMM as continuous visual embeddings, which cannot obtain discrete supervised labels for classification. As a compromise solution, mainstream LMMs choose to compute losses only for the language portion of multi-modal interleaved sequence. As illustrated in figure  2(a), the training objective of visual context method (Dai et al., 2023; Li et al., 2023b; Liu et al., 2023b; Lu et al., 2023; Bai et al., 2023a) focuses on predicting language responses in multi-modal contextual sequences that depend on visual information, where the visual information merely acts as contextual cues and does not serve as supervision. This unfair treatment of different modal information in multi-modal sequences lacks the process of learning different modal information utilising the inference capabilities of the LLM, severely limiting the potential of LMM and resulting in under-utilisation of the training data. Although recent works  (Sun et al., 2023a, b) propose to unlock the LLM for text pre-training by using a regression task to predict the value of next visual feature in the pre-training phase (Fig.  2(b)), the inconsistent optimisation goals of its visual and linguistic components are not conducive to unified multi-modal auto-regressive modelling.  (Bai et al., 2023b) has also proposed learning visual features using an auto-regressive classification tasks, but it uses a pre-trained image tokenizer, such as VQVAE or VQGAN, to cluster image features into a grid of discrete tokens which does not combine discrete visually supervised information with LLM.
@@
In this paper, we successfully perform multi-modal auto-regressive modeling with a unified objective for the first time. Specifically, we introduce the concept of visual tokens to construct representations of visual features in the language semantic space inside the LMM, thus implementing visual modelling of LMM in the form of classification rather than regression. We also propose the Visual Token guided Large Multi-modal Model (VT-LMM), a novel multi-modal model that inherits the successful learning paradigm of LLM in the pre-training task, i.e., predicting the next image/text token in an auto-regressive manner. With the help of visual tokens, it’s possible to train LMM with auto-regressive modeling without any specific architectural modifications, as shown in figure  2(c). Just as the LM head component is used in LLM to accomplish the mapping of text features to interpretable semantics, we correspondingly construct the VM head component, using the same structure, intent on obtaining the representation of each visual embedding in the language semantic space of LLM. For each visual embedding, we use the VM head to map it into probability distributions over the pre-trained vocabulary, which we call visual tokens. After that, the visual tokens corresponding to the image modal and the one hot label corresponding to the text modal are intertwined to form the supervised information for multi-modal auto-regressive modelling. Further, the pre-trained embeddings weights of LLM can be approximated as a set of complete bases of the language semantic space, i.e., the semantic information covered by the embeddings can basically cover the whole semantic space. Therefore, we further explored whether pseudo image features constructed with visual tokens and LLM’s pre-trained embeddings can convey visual information to the model. We conducted extensive experiments on five commonly used visual question answering benchmarks and four LMM-evaluating benchmark toolkits, and the experimental results demonstrate that our VT-LMM, by constructing visual tokens to introduce visual supervisory information, achieves the best performance among models of the same scale, and obtains vision-language understanding capability competitive to or even surpassing that of 13B or even larger scale models with a scale of 7B. Our main contributions are as follows:
@@
  • We propose the concept of visual tokens, which maps visual features into language semantic space, enabling LMM to perform auto-regressive modelling over multi-modal sequence.
@@
2. Method
Figure 3. The overview of our method. (a) The overall framework of the model. VT-LMM uses the VM head to transform visual features in multi-modal input sequences to probability distributions over LLM’s vocabulary (so-called visual tokens) to participate in visual modelling (b) Constructing pseudo image features with pre-trained embedding of LLM and visual tokens. (c) Demonstration of semantically closest tokens of each image patch in LMM.
2.1. Multi-modal Learning
To extend the powerful text-only perceptual and reasoning capabilities of LLM to scenarios dealing with multi-modal inputs, existing multi-modal learning method of LMM typically use the adapter structure to transform the visual features encoded by pre-trained visual backbone into the semantic space of LLM and construct multi-modal input sequences together with text embeddings. The adapter could be well-designed visual resampler (Bai et al., 2023a; Dai et al., 2023), Multi-layer Perceptron (MLP) (Liu et al., 2023b) or even simple Linear layer (Liu et al., 2023c).
@@ 2.1. Multi-modal Learning
(1) $X_{\mathrm{image}}=\mathrm{AD}(\mathrm{VE}(\mathrm{image})),$
@@
Subsequently, $X_{\mathrm{input}}$ is fed into the LLM for auto-regressive decoding. Assuming that the LLM mainly contains two important components: a language decoder $\Gamma$ and an LM head, the loss of conventional multi-modal learning can be expressed as
(2) $Loss_{\mathrm{LM}}^{\mathrm{Conv}}$ $=\frac{1}{|S_{\mathrm{{LM}}}|}\sum_{n\in S_{\mathrm{{LM}}}}\left(-\sum_{i=0}^{C-1}P_{n}(i)\log Q_{n}(i)\right),$
@@
where $W_{\mathrm{LM}}$ is the optimizable parameters of $\mathrm{LM\ head}$, set $S_{\mathrm{{LM}}}$ is the set of expected text output index of model, $P$ represents the ground truth label and $C$ is the vocabulary size of model.
However, classical multi-modal learning neglects the supervision function of visual information on multi-modal auto-regressive modelling, which is still multi-modal based language modelling in essence. Our work proposes to leverage the great reasoning potential of LLM to facilitate both language modelling and visual modelling of LMM. To achieve this goal, the key component is to construct formally unified multi-modal generative objectives so that the model can be trained with auto-regressive manner. As shown in Figure 3(a), the architecture of VT-LMM consists of five components: a Visual encoder, a multi-modal decoder, an adapter for projection between visual modal and language modal, a Multi-modal Modeling head (MM head) for multi-modal modelling and a corresponding VM head for visual modelling. The embedding module can be regarded as part of multi-modal decoder. We initialise the multi-modal decoder and MM head in VT-LMM using the pre-trained LLM and its LM head. In contrast to classical multi-modal learning, VT-LMM uses VMhead to construct visual tokens for visual features as supervisory information, thus enabling the model to perform multi-modal auto-regressive modeling over the entire sequence.
2.2. Visual Tokens
@@
(3) $t_{i}=\arg\min_{j}\operatorname{Cosine}(x_{i},e_{j}),\quad t_{i}\in[0,C-1]$
@@
We show the semantically closest tokens of image regions in figure 3(c), it can be observed that for the LMM with multi-modal alignment training, there is a certain correlation between the information contained in each vector of the image features and the actual content of the image. Moreover, this correlation can be explicitly expressed by the vocabulary of the LMM. This finding, on the one hand, proves that the visual features within the LMM are in a similar semantic space as the text embedding, and on the other hand, it provides a new perspective for the interpretability analysis of the LMM. Furthermore, existing LLM methods have shown that the hidden states of text features can be mapped to the model’s vocabulary through linear projection, enabling the extraction of interpretable semantics. Therefore, we propose using the linear projection to also map the visual features to the probability distribution over the model’s vocabulary, which we called visual tokens, to further strengthen the correlation between visual features and text embeddings.
@@
(4) $Loss_{\mathrm{LM}}$ $=\frac{1}{|S_{\mathrm{{LM}}}|}\sum_{n\in S_{\mathrm{{LM}}}}\left(-\sum_{i=0}^{C-1}P_{n}(i)\log Q_{n}(i)\right),$
@@
The visual tokens of visual features can be represented as
(5) $P^{\prime}$ $=\operatorname{Softmax}(W_{\mathrm{VM}}X_{\mathrm{image}}),$
@@
(6) $Loss_{\mathrm{VM}}$ $=\frac{1}{|S_{\mathrm{{VM}}}|}\sum_{n\in S_{\mathrm{{VM}}}}\left(\sum_{i=0}^{C-1}P^{\prime}_{n}(i)\log(\frac{P^{\prime}_{n}(i)}{Q_{n}(i)})\right).$
where set $S_{\mathrm{{VM}}}$ is the set of visual output index of model. It should be noted that in the training process of VT-LMM, the optimisation phase of VM head and multi-modal decoder are separated, and thus do not cause instability during training.
@@
(7) $Loss_{\mathrm{MM}}=Loss_{\mathrm{LM}}+Loss_{\mathrm{VM}}.$
@@
2.3. Fused Text Embeddings as Pseudo Image Features
The text embeddings in the multi-modal decoder can be regarded as a set of base vectors within its semantic space. Therefore, we propose utilizing the visual tokens and text embeddings to construct pseudo image features, aiming to further explore the manifestation of visual features in the semantic space of the LMM.
@@ 2.3. Fused Text Embeddings as Pseudo Image Features
(8) $X_{\mathrm{p}}=W_{\mathrm{embeddings}}\odot\operatorname{Softmax}\left((W_{\mathrm{VM}}X_{\mathrm{image}})\right)$
@@ 3. Experiments
IDEFICS-80B (IDEFICS, 2023) LLaMA-65B 224 60.0 45.2 36.0 30.9 54.5 38.1
InstructBLIP (Dai et al., 2023) Vicuna-13B 224 49.5 33.4 63.1 50.7 78.9 25.6
BLIP-2 (Li et al., 2023b) Vicuna-13B 224 41.0 41.0 19.6 61.0 42.5 85.3 22.4
LLaVA-v1.5 (Liu et al., 2023b) Vicuna-13B 336 80.0 63.3 53.6 71.6 61.3 85.9 67.7 63.6 35.4
InstructBLIP (Dai et al., 2023) Vicuna-7B 224 49.2 34.5 60.5 50.1 36 23.7 26.2
IDEFICS-9B (IDEFICS, 2023) LLaMA-7B 224 50.9 38.4 35.5 25.9 48.2 25.2
Qwen-VL (Bai et al., 2023a) Qwen-7B 448 78.8 59.3 35.2 67.1 63.8 38.2 7.4
Qwen-VL-Chat (Bai et al., 2023a) Qwen-7B 448 78.2 57.5 38.9 68.2 61.5 60.6 56.7
LLaVA-v1.5 (Liu et al., 2023b) Vicuna-7B 336 78.5 62.0 50.0 66.8 58.2 85.9 64.3 58.3 30.5
MoE-LLaVA-2.7B×4-Top2 (Lin et al., 2024) Phi-2-2.7B 336 77.6 61.4 43.9 68.5 51.4 86.3 65.2 34.3
@@
Emu2-Chat (Sun et al., 2023a) LLaMA-33B 448 84.9 65.1 54.9 65.5 66.6 48.5
Emu-I (Sun et al., 2023b) LLaMA-13B 224 62.0 46.0 38.3 36.3
MM-Interleaved-SFT (Tian et al., 2024) Vicuna-13B 224 80.2 60.5 54.9 61.0
Unified-IO 2 (Lu et al., 2023) UIO-2-6.8B 384 79.4 86.2 87.7 71.5
DreamLLM (Dong et al., 2023) Vicuna-7B 224 56.6 38.1 34.9
VL-GPT-I (Zhu et al., 2023) LLaMA-7B 224 67.2 51.5 38.9
LaVIT-v2 (Jin et al., 2023) LLaMA2-7B 224 68.3 47.9 41.0
VT-LMM Vicuna-7B 336 78.9 62.7 48.3 68.1 57.6 85.9 65.9 59.8 31.3
VT-LMM Mistral-7B 336 80.8 65.4 58.5 75.9 63.1 87.0 80.6 79.0 44.0